lizzy69:

On vacation, Li’s family will be visiting an amusement park. Riders must be at least 48 inches tall to go on the roller coaster. Li’s little brother is 3 feet 6 inches tall. Which inequality can be used to find how many inches he must grow to be able to ride the roller coaster? (12 in. = 1 ft)

anthony5012005:

so the rider has to be 48in and little brother is 42 in and u r looking for h so thats why it is 48<__h+42

Letsgoooo:

the little brother must grow 6 inches to ride the roller coaster. 12 (in) x 3 + 6= 42, therefore 42+ 6= 48 (in). I hope this help u